OpenStudy (mathmath333):

The probability of bomb hitting a bridge is 1/2 and two direct hits are required to destroy it.The least number of bombs required so that the probability of the bridge being destroyed is greater than 0.9 is

OpenStudy (mathmath333):

\(\large \color{black}{\begin{align} & \normalsize \text{The probability of bomb hitting a bridge is}\ \dfrac12\ \text{and two direct hits} \hspace{.33em}\\~\\ & \normalsize \text{ are required to destroy it.The least number of bombs}\hspace{.33em}\\~\\ & \normalsize \text{ required so that the probability of the bridge being destroyed}\hspace{.33em}\\~\\ & \normalsize \text{ is greater than 0.9 is }\hspace{.33em}\\~\\ & a.)\ 7\ \ bombs \hspace{.33em}\\~\\ & b.)\ 3\ \ bombs \hspace{.33em}\\~\\ & c.)\ 8\ \ bombs \hspace{.33em}\\~\\ & d.)\ 9\ \ bombs \hspace{.33em}\\~\\ \end{align}}\)

OpenStudy (phi):

is this a poisson distribution ?

OpenStudy (mathmath333):

i found a page i cant understand soln here http://mathhelpforum.com/statistics/81015-probability-problems.html

OpenStudy (phi):

ok, or binomial

OpenStudy (mathmath333):

i hvent studied poision ditributuion , i have this in probability chapter

OpenStudy (phi):

then binomial looks appropriate

OpenStudy (mathmath333):

i also didnt studied binomial distributuion

OpenStudy (phi):

what techniques are in this chapter ?

OpenStudy (mathmath333):

Use of AND use of OR Combination of AND and OR Random experiment sample space Event Non-event Imposiible event mutually exclusive event independent event conditional probablity

OpenStudy (phi):

I guess we can use all of that to derive the binomial distribution, and then solve the problem. For example, what is the chances we have 0 hits in n tries ? Can you do that problem ?

OpenStudy (mathmath333):

1/2

OpenStudy (phi):

it is 1/2 if we tried once but what if we try n times ?

OpenStudy (mathmath333):

(1/2)^n

OpenStudy (phi):

yes. next, what are the chances we get exactly 1 hit in n tries ?

OpenStudy (mathmath333):

(1/2)^n

OpenStudy (phi):

First, what is the prob of hitting on the first shot, and missing on all the others ? 1/2 * 1/2^(n-1) = 1/2^n but we could miss the first shot, hit on the second, and miss on the rest. that is also 1/2^n and so on...

OpenStudy (phi):

in other words, the chance of making exactly one hit is n * (1/2)^n ok ?

OpenStudy (mathmath333):

ok

OpenStudy (phi):

so we know the chances of getting exactly 0 or 1 hit 1 - that prob is the chance of getting 2 (or more) hits in n tries.

OpenStudy (phi):

in other words, Pr(2 or more hits) = \( 1 - \frac{1}{2^n} - \frac{n}{2^n} \) we want that to be bigger than 0.9 this is not easy to solve, except numerically, so use trial and error.

OpenStudy (mathmath333):

in other words, the chance of making exactly one hit is \(\color{red}{n}\) * (1/2)^n ok ? where does \(\color{red}{n}\) comes from

OpenStudy (phi):

say we try 3 times then we could have x 0 0 0 x 0 0 0 x each line has a 1/8 chance and we have a total of 3/8 chance of getting exactly one hit.

OpenStudy (mathmath333):

thnks